# SMSupplementary References

### v1

@@ Abstract
Are vision-language models (VLMs) open-set models because they are trained on internet-scale datasets? We answer this question with a clear no – VLMs introduce closed-set assumptions via their finite query set, making them vulnerable to open-set conditions. We systematically evaluate VLMs for open-set recognition and find they frequently misclassify objects not contained in their query set, leading to alarmingly low precision when tuned for high recall and vice versa. We show that naively increasing the size of the query set to contain more and more classes does not mitigate this problem, but instead causes diminishing task performance and open-set performance. We establish a revised definition of the open-set problem for the age of VLMs, define a new benchmark and evaluation protocol to facilitate standardised evaluation and research in this important area, and evaluate promising baseline approaches based on predictive uncertainty and dedicated negative embeddings on a range of VLM classifiers and object detectors.
@@ 1 Introduction
In 2013, Scheirer et al. [40] described the closed-set assumption ingrained in almost all vision models at the time: all test classes are known during training and the model is never exposed to novel or unknown classes during testing. Real-world applications can challenge this assumption and impose open-set conditions – unexpected objects not included in the training dataset – causing potentially dangerous overconfident misclassifications [11, 41, 28, 4]. Open-set recognition [40] was introduced to address this issue by evaluating the ability of vision models to identify and reject open-set inputs as unknown.
Vision-language models (VLMs) have recently revolutionised the fields of image classification [35, 21, 16] and object detection [51, 52, 55, 17, 56, 24, 47]. Trained on vast and diverse internet-scale datasets, VLMs recognise an extensive variety of classes and seemingly are open-set by default [45]. Our paper challenges this overly optimistic view.
In the age of VLMs, we introduce an updated definition of the open-set problem – while previous definitions focused on a finite training set [40], we show that VLMs impose closed-set assumptions through a finite query set. As illustrated in Fig. 1, VLMs are currently still evaluated under closed-set conditions: VLMs compare image embeddings with text embeddings from a predefined query set of class labels. This introduces a closed-set assumption that all classes encountered during testing are included in this predefined query set. Open-set conditions emerge when VLMs encounter objects that are not included in this predefined query set (unknown objects). As we will show in Section 5, even state-of-the-art VLMs heavily degrade in performance and misclassify unknown objects as belonging to the query set with high confidence.
Figure 1: Left: Traditional open-set recognition arises when testing classes outside a model’s finite training set. While VLMs are trained on internet-scale datasets containing most conceivable object classes, VLMs still use a finite query set for classification. Open-set recognition arises when test classes are not included in the query set. Right: When VLMs are tested with an object class included in the predefined query set, they often perform well and correctly classify the object. When tested with an object class that is not present in the query set (i.e. testing an open-set object), VLMs often misclassify the object as a query class with high confidence (i.e. an open-set error).
@@ 2.1 Vision-Language Models for Zero-shot Classification
Recent years have seen the emergence of foundation models for computer vision. Trained on internet-scale datasets with self-supervision learning techniques, foundation models can be applied to a range of downstream tasks [6]. One type of foundation model is the vision-language model (VLM), which learns a multi-modal feature space that can jointly represent visual features and text [35, 21, 49, 53, 1] (and in some cases other modalities as well [50, 16]).
Large-scale learning provides VLMs with a rich feature representation that can be adapted for many downstream tasks, one of which is zero-shot or open vocabulary image classification [35, 21, 49, 53, 1, 50, 16].
Despite not being trained for classification tasks or datasets, VLMs generalise surprisingly well – for example, VLM CoCa [49] is within $5\%$ of the current supervised state-of-the-art for top-1 accuracy on ImageNet. Alongside this, one of the distinguishing advantages of VLMs for classification is their adaptability to different datasets with different classes – the set of classes can be changed at will depending on the dataset [35, 21].
@@ 2.2 Vision-Language Models for Open Vocabulary Object Detection
Unlocked by the advances in VLMs for zero-shot classification, [52] proposed the task of open vocabulary object detection. Traditionally, object detection involves training with an exhaustively labelled dataset to detect a fixed set of object classes [13, 25]. In contrast, open vocabulary object detection requires detectors that can generalise to an arbitrary set of object classes at test time, even object classes that were unknown during the training of the detector [52].
Existing open vocabulary detectors leverage VLM backbones to enable generalisation and adaptability to new object classes [52, 17, 51, 55, 56, 24, 47, 12, 22, 46, 14, 33]. These detectors are generally inspired by a Faster R-CNN-style architecture (with a region proposal network) [52, 17, 56, 55, 24, 12, 22, 46, 14, 33, 44] or DETR-style architecture [51, 47]. CLIP [35] is frequently utilised by these open vocabulary detectors, directly included in the architecture via its text encoder [55, 51, 17, 24, 56, 12, 46, 14, 33, 44] or in its entirety [47, 22], or used as a supervision signal during training  [55, 17, 51, 12, 46, 14, 33, 44].
We refer the reader to [45] for a comprehensive survey on open vocabulary learning.
@@ 2.3 Open-set Recognition
Scheirer et al. [40] introduced open-set recognition to challenge the prevailing closed-set assumptions of image recognition systems. Closed-set models only test on samples from a predefined set of classes that were known during model training [40]. In open-set recognition, samples can instead come from an expansive and ever-changing range of classes that cannot be predefined, and therefore may be unknown during model training [40]. The goal of open-set recognition is to equip models with the ability to identify and reject unknown samples that do not belong to one of the training classes. The open-set concept has been explored widely in computer vision, including in image classification [4, 43, 54, 48, 30, 32, 8], object detection [11, 28, 18, 29, 57], image segmentation [19, 34, 7] and other vision tasks [31, 38, 27, 3, 26].
@@ 3.1 Baseline Approaches to Open-Set Recognition for VLMs
We investigate a number of baseline options for negative embeddings: a simple word approach, where we use “This is a photo.” without specifying a class label; a random words method, where we generate random strings with random length between 2-8 characters (e.g. “braxl”); a zero embedding approach, with an embedding vector of all zeroes, i.e. $\mathcal{E}^{\texttt{n}}=\{\mathbf{0}\}$; and a random embeddings method, where we randomly draw $M$ embedding vectors from a Gaussian distribution, i.e. $\mathbf{e}^{\texttt{n}}_{i}\sim\mathcal{N}({\boldsymbol{\mu}},{\boldsymbol{\Sigma}})$, using the mean and standard deviation from the query embedding vectors $\mathcal{E}^{q}$. We compare the effectiveness of these approaches, including for different numbers of $M$ random negative queries, in our experiments. Interestingly, many open vocabulary object detectors already use the zero embedding approach as a negative query approach [52, 55, 56, 24, 47].
@@ 4 Evaluation Protocol
Our evaluation protocol is publicly available at this link (inserted upon paper acceptance), along with the files to reproduce all results presented in this paper.
@@ 4.1 Creating an Open-set Recognition Dataset for VLMs
Testing Closed-set Recognition: We first test images in a standard closed-set manner to establish the closed-set performance of the VLM and identify all true positive (TP) predictions. This involves testing all images with a query set identical to the dataset class list, $\mathcal{Q}=\mathcal{L}$.
To identify the TP predictions, we follow standard protocol for classification and object detection datasets – in classification, a TP has a predicted class matching the ground-truth label; in object detection, this involves a object-detection assignment process that requires correct object classification and localisation. We include more details in the Supplementary Materials. TPs represent correct closed-set predictions that should not be rejected as unknown, and we collect the uncertainty measures $\psi$ associated with all TPs for later evaluation.
@@ 4.2 Metrics
We use performance metrics that measure the ability of the VLM to reject OSEs (with high uncertainty or with a negative class), while maintaining the closed-set (task) performance.
Precision-Recall Curves: Open-set recognition can be formulated as a binary classification task that uses uncertainty as the decision threshold. Correct closed-set predictions (introduced as TPs in Sec. 4.1) and open-set errors (OSEs) are considered as the positive and negative class respectively. We construct precision-recall curves, which are well-suited to this problem as they are affected by the size of the negative class (OSEs) [39] [^4] ROC curves do not have this property – we discuss this further in the Supplementary Material where we also report AuROC for all experiments. as well as prediction uncertainty. High precision indicates that little-to-none negatives (OSEs) remain after the uncertainty thresholding. High recall indicates that most positives (TPs) remain after the uncertainty thresholding. We summarise the performance of the precision-recall curves with 3 metrics: (1) Area under the PR curve (AuPR), (2) Precision at $\mathbf{95\%}$ Recall, and (3) Recall at $\mathbf{95\%}$ Precision.
@@ 4.3 VLM Baselines
We test with three VLM classifiers – namely CLIP [35], ALIGN [21] and ImageBind [16] – and seven open vocabulary object detectors – namely OVR-CNN [52], ViLD [17], OV-DETR [51], RegionCLIP [55], Detic [56], VLDet [24] and CORA [47]. These VLMs cover a range of architectures and training paradigms and have publicly-available implementations available. Our goal with testing this range of VLMs is not to find the “best” VLM for open-set recognition, but instead to exhibit the general vulnerability of VLMs to open-set. For each VLM, we extract the prediction Softmax score, Cosine similarity, and Entropy of the softmax distribution as baseline uncertainty measures (see Sec. 3.1). Some open vocabulary object detectors use Sigmoid activation rather than Softmax [51, 56, 24, 47] – for these methods, we only report the Sigmoid score as its results are identical to the cosine similarity (one-to-one mapping) and the entropy is unsuitable as the scores are not normalised to a distribution. Additional implementation details for each model are included in the Supplementary Material.
@@ 4.4 Datasets
We focus on general object recognition and test with the large-scale and prevailing datasets used to benchmark image classification and object detection: For image classification, we test with the ImageNet1k validation dataset [10, 37]. It contains $50,000$ images from $1000$ classes and is publicly available for research. For object detection, we test with the MS COCO 2017 validation dataset [25]. It contains $4,952$ images with $36,781$ annotated objects from $80$ classes and is publicly available for research. We show results for domain-specific datasets in the Supplementary Material.
@@ 5 Experiments and Results
AuPR P@95R R@95P TP OSE Accuracy
Classifier Uncertainty $\uparrow$ $\uparrow$ $\uparrow$ $\uparrow$ $\downarrow$ $\uparrow$
CLIP [35] Softmax 71.5 46.2 3.4 31,023 50,000 62.1
Cosine 61.3 42.5 0.1
Entropy 72.3 45.4 2.8
ALIGN [21] Softmax 74.7 48.1 8.3 32,618 50,000 65.2
Cosine 62.1 44.8 -
Entropy 75.0 46.8 8.6
ImageBind [16] Softmax 79.1 52.8 5.0 38,405 50,000 76.8
Cosine 73.9 50.7 0.2
Entropy 80.1 56.3 5.6
Tab. 1 shows that all three VLM classifiers, CLIP [35], ALIGN [21] and ImageBind [16], perform poorly for open-set recognition. For scenarios that require high recall of true positive predictions (i.e. $95\%$ recall), approximately every second prediction returned is an open-set error (i.e. precision between $46.2\%$ and $56.2\%$). For scenarios that require high precision of predictions (i.e. $95\%$ precision), less than 10% of the true positive predictions are retained (i.e. recall between 3.4% and 8.6%). Comparing across uncertainty types, all VLM classifiers show better uncertainty performance when using the predicted Softmax score or Entropy rather than the predicted Cosine similarity. This suggests raw similarity in VLM feature space is not the best indicator of open-set error, with measures of relative similarity compared to other classes offering better performance.
@@
ViLD (R152)[17] Softmax “background" 60.1 7.5 15.4 16,011 1,485,600 46.4
Cosine “background" 33.1 6.9 0.1
Entropy “background" 62.2 8.4 16.3
@@
High open-set error counts are not inherently a problem if they are always produced with high uncertainty. Yet for all tested detectors, the baseline uncertainty methods cannot adequately distinguish true positive detections from open-set errors. When thresholding for high recall of true positives, at best only every fifth detection is a true positive (i.e. precision ranges between $5.9\%$ and $21.9\%$). When thresholding for high precision, at best less than half of true positives are recalled (i.e. recall ranges between 16.3% and 48.5%).
Notably, the precision-recall characteristics for VLM classification and object detection differ significantly – VLM classifiers achieve better performance at high recall, whereas VLM detectors achieve better performance at high precision. While the VLM classifiers struggle with over-confident open-set errors, the VLM detectors suffer from under-confident true positives. We include histograms that visualise this trend in the Supplementary, as well as full precision-recall curves.
@@ 5.2 Can Negative Queries Improve Open-set Performance?
Table 3: The open-set performance of state-of-the-art VLM classifiers when combined with negative embeddings, tested on ImageNet1k. Arrows indicate direction of optimal performance. Negative embeddings introduced as words are higlighted gray.
@@
Simple-W 71.5 46.2 3.4 31,023 49,982 62.0
CLIP [35] 500 Rand-W 73.4 48.1 3.3 29,961 41,428 59.9
(Softmax) 2500 Rand-W 72.8 49.8 1.6 28,657 35,092 57.3
Zero-Emb 71.5 46.2 3.4 31,019 49,955 62.0
500 Rand-Emb 71.9 47.2 2.5 30,160 44,673 60.3
2500 Rand-Emb 71.7 47.1 0.9 29,291 40,497 58.6
@@
Simple-W 74.7 48.1 8.2 32,604 49,832 65.2
ALIGN [21] 500 Rand-W 75.3 48.9 10.4 31,731 43,198 63.5
(Softmax) 2500 Rand-W 74.6 49.5 8.9 30,588 37,738 61.2
Zero-Emb 74.7 48.1 8.3 32,618 50,000 65.2
500 Rand-Emb 74.9 48.7 8.6 32,342 47,228 64.7
2500 Rand-Emb 75.1 49.7 7.2 31,462 41,210 62.9
@@
Simple-W 79.3 53.2 5.7 38,343 49,008 76.7
ImageBind [16] 500 Rand-W 83.8 63.1 7.7 35,859 29,172 71.7
(Softmax) 2500 Rand-W 84.6 66.1 8.3 34,179 22,872 68.4
Zero-Emb 79.1 52.8 5.2 38,405 49,998 76.8
500 Rand-Emb 81.0 57.3 9.0 36,966 38,636 73.9
2500 Rand-Emb 81.7 59.0 11.2 35,954 33,349 71.9
Figure 2: The trade-off between closed-set and open-set performance with increasing negative queries and VLM classifiers.
Classification: Negative class queries provide the VLM classifiers with the ability to select “unknown” when testing an input. When used in larger quantities, negative queries are able to capture and remove significant portions of the open-set error, yet this happens in trade-off with a reduction in closed-set performance – see Tab. 3. See for example ImageBind, where using $2,500$ random word negative queries effectively halves the open-set error (i.e. more than $25,000$ open-set images are classified as one of the random words), but this occurs in trade-off with a closed-set accuracy reduction of $8.4\%$ (i.e. approx. $4,000$ of the closed-set images are incorrectly classified as one of the random words). We present the rate of this trade-off against the number of negative query points for the different classifiers and negative query types (word or embedding) in Fig. 2. For all VLM classifiers, negative queries introduced as random words appear to more rapidly reduce both open-set error and accuracy.
While negative queries can noticeably reduce the raw open-set error numbers, this is not always mirrored with increased performance in the open-set uncertainty metrics (i.e. AuPR). In some cases, a decrease in the uncertainty performance is even observed. This indicates that negative queries may be capturing and removing the “easy” open-set errors, i.e. the open-set errors that already had high uncertainty – this is confirmed in Fig. 3, which shows which predictions are removed when introducing $500$ random words as negative queries. In contrast to CLIP and ALIGN, ImageBind is able to remove previously overconfident open-set errors with these negative queries, resulting in up to $5.5\%$ increase in AuPR.
Figure 3: Using negative queries (random 500 words) removes open-set error with different uncertainties depending on the VLM classifier.
Object Detection: The VLM detectors do not show a consistent response to the use of negative queries. Different detectors show different sensitivities to negative query counts, different trade-off rates between closed-set and open-set performance, and differences in effectiveness of embedding versus word negative query types – see Tab. 4 and Fig. 4.
@@
Looking at Fig. 4, we can consider the sensitivity, trade-off rate and response of the detectors to negative queries: greater performance changes with increasing queries indicates sensitivity; large decreases in mAP with more queries indicates increased trade-off rates between open-set and closed-set performance; and a positive response to the negative queries presents as an increase in AuPR when increasing queries (negative response decreases AuPR).
Figure 4: The trade-off between closed-set and open-set performance with increasing negative queries and VLM object detectors.

### v2

@@ Abstract
Are vision-language models (VLMs) for open-vocabulary perception inherently open-set models because they are trained on internet-scale datasets? We answer this question with a clear no – VLMs introduce closed-set assumptions via their finite query set, making them vulnerable to open-set conditions. We systematically evaluate VLMs for open-set recognition and find they frequently misclassify objects not contained in their query set, leading to alarmingly low precision when tuned for high recall and vice versa. We show that naively increasing the size of the query set to contain more and more classes does not mitigate this problem, but instead causes diminishing task performance and open-set performance. We establish a revised definition of the open-set problem for the age of VLMs, define a new benchmark and evaluation protocol to facilitate standardised evaluation and research in this important area, and evaluate promising baseline approaches based on predictive uncertainty and dedicated negative embeddings on a range of open-vocabulary VLM classifiers and object detectors.
@@ 1 Introduction
In 2013, Scheirer et al. [38] described the closed-set assumption ingrained in almost all vision models at the time: all test classes are known during training and the model is never exposed to novel or unknown classes during testing. Real-world applications challenge this assumption and test in open-set conditions – encountering unexpected objects not in the training dataset – and causing potentially dangerous overconfident misclassifications [10, 39, 27, 4]. Open-set recognition [38] was introduced to address this issue by evaluating the ability of vision models to identify and reject open-set inputs as unknown.
Vision-language models (VLMs) have recently revolutionised the fields of image classification [34, 19, 14] and object detection [49, 51, 15]. Trained on vast and diverse internet-scale datasets, VLMs recognise an extensive variety of classes and seemingly are open-set by default [43]. Our paper challenges this view.
In the age of VLMs, we introduce an updated definition of the open-set problem – while previous definitions focused on a finite training set [38], we show that VLMs impose closed-set assumptions through a finite query set. As illustrated in Fig. 1, VLMs for open-vocabulary image classification or object detection are still evaluated under closed-set conditions: they compare image embeddings with text embeddings from a predefined, finite query set of class labels. This introduces a closed-set assumption that all classes encountered during testing are included in this query set. Open-set conditions emerge when VLMs encounter objects that are not included in the query set (unknown objects). As we show in Section 5, even state-of-the-art VLMs heavily degrade in performance and misclassify unknown objects as belonging to the query set with high confidence.
Figure 1: Left: Traditional open-set recognition arises when testing classes outside a model’s finite training class set. While VLMs are trained on internet-scale datasets containing most conceivable object classes, they use a finite query set for classification. Open-set recognition arises when test classes are not included in the query set. Right: When testing an object present in the predefined query set, VLMs often correctly classify the object. When testing an object that is not present in the query set (i.e. an open-set object), VLMs often misclassify the object as a query class with high confidence (i.e. an open-set error).
@@ 2.1 Vision-Language Models for Zero-shot Classification
A recently emerged foundation model is the vision-language model (VLM), which learns a multi-modal feature space that can jointly represent visual features and text [34, 19, 47, 52, 1] (and in some cases other modalities as well [48, 14]). Large-scale contrastive, self-supervised learning on internet-scale datasets provides VLMs with a rich feature representation that can be adapted for many downstream tasks, one of which is zero-shot or open vocabulary image classification [34, 19, 47, 52, 1, 48, 14]. Despite not being trained for classification tasks or datasets, VLMs generalise surprisingly well and achieve near state-of-the-art supervised learning performance [47]. Alongside this, a distinguishing advantage of VLMs is their adaptability to different datasets with different classes – the set of classes can be changed at will depending on the dataset [34, 19].
@@ 2.2 Vision-Language Models for Open Vocabulary Object Detection
Unlocked by the advances in VLMs for zero-shot classification, [51] proposed open vocabulary object detection, where detectors must generalise to an arbitrary set of object classes at test time, even object classes that were not seen during training [51]. This is related to our proposed open-set recognition, but assumes that novel object classes are added to the query set, whereas we test against object classes that are not present in the query set.
Existing open vocabulary detectors leverage VLM backbones to enable generalisation to new object classes, embedded within Faster R-CNN-style architectures [51, 15, 56, 55, 22, 11, 20, 44, 12, 32, 42] or DETR-style architectures [49, 45]. CLIP [34] is frequently utilised by these open vocabulary detectors, directly included in the architecture via its text encoder [55, 49, 15, 22, 56, 11, 44, 12, 32, 42] or in its entirety [45, 20], or used as a supervision signal during training  [55, 15, 49, 11, 44, 12, 32, 42]. We refer the reader to [43] for a comprehensive survey on open vocabulary learning.
@@ 2.3 Open-set Recognition
Scheirer et al. [38] introduced open-set recognition to challenge the prevailing closed-set assumptions of image recognition systems. Closed-set models only test on samples from a predefined set of classes that were known during model training [38]. In open-set recognition, samples can instead come from an expansive range of classes that cannot be predefined, and therefore are unknown during model training [38].Open-set recognition tests the ability of models to identify and reject unknown samples that do not belong to one of the training classes. The open-set concept has been explored widely in computer vision, including in image classification [4, 41, 54, 46, 29, 31, 7], object detection [10, 27, 16, 28, 57], image segmentation [17, 33, 6] and other vision tasks [30, 36, 26, 3, 25].
@@ 3.1 Baseline Approaches to Open-Set Recognition for VLMs
We investigate two baseline options for negative embeddings: a random words method, where we generate random strings with random length between 2-8 characters (e.g. “braxl”); and a random embeddings method, where we randomly draw $M$ embedding vectors from a Gaussian distribution, i.e. $\mathbf{e}^{\texttt{n}}_{i}\sim\mathcal{N}({\boldsymbol{\mu}},{\boldsymbol{\Sigma}})$, using the mean and standard deviation from the query embedding vectors $\mathcal{E}^{q}$. We compare the effectiveness of these approaches, including for different numbers of $M$ random negative queries, in our experiments. Interestingly, many open vocabulary object detectors already use a negative query approach, though typically with a only single embedding vector of all zeroes [51, 55, 56, 22, 45].
@@ 4.1 Creating an Open-set Recognition Dataset for VLMs
Testing Closed-set Recognition: We first test images in a standard manner to establish the closed-set performance of the VLM and identify all true positive (TP) predictions. This involves testing all images with a query set identical to the dataset class list, $\mathcal{Q}=\mathcal{L}$. To identify the TP predictions, we follow standard protocol for classification and object detection datasets – in classification, a TP has a predicted class matching the ground-truth label; in object detection, this involves a object-detection assignment process that requires correct object classification and localisation (see Supplementary Sec. S1). TPs represent correct closed-set predictions that should not be rejected as unknown, and we collect the uncertainty measures $\psi$ associated with all TPs for later evaluation.
@@ 4.2 Metrics
We use performance metrics that measure the ability of the VLM to reject OSEs (with high uncertainty or with a negative class), while maintaining the closed-set task performance.
Precision-Recall Curves: Open-set recognition can be formulated as a binary classification task that uses uncertainty as the decision threshold. Correct closed-set predictions (introduced as TPs in Sec. 4.1) and open-set errors (OSEs) are considered as the positive and negative class respectively. We construct precision-recall curves, which are well-suited to this problem as they are affected by the size of the negative class (OSEs) [37] [^4] ROC curves do not have this property – we discuss this further in Sec. S3 of the Supplementary Material where we report AuROC for the object detection experiments. as well as prediction uncertainty. High precision indicates that little-to-none negatives (OSEs) remain after the uncertainty thresholding. High recall indicates that most positives (TPs) remain after the uncertainty thresholding. We summarise the performance of the precision-recall curves with 3 core metrics: (1) Area under the PR curve (AuPR), (2) Precision at $\mathbf{95\%}$ Recall, and (3) Recall at $\mathbf{95\%}$ Precision.
@@ 4.3 VLM Baselines
We test with six VLM classifiers – CLIP [34], ALIGN [19], CoCa [47], ImageBind [14], SigLIP [53] and LanguageBind [59] – and seven open vocabulary object detectors – OVR-CNN [51], ViLD [15], OV-DETR [49], RegionCLIP [55], Detic [56], VLDet [22] and CORA [45] (see Supplementary Sec. S2 for implementation details). These VLMs cover a range of architectures and training paradigms and have publicly-available implementations. Our goal with testing this range of VLMs is not to find the “best” VLM for open-set recognition, but instead to exhibit the general vulnerability of VLMs to open-set conditions. For each VLM, we extract the prediction Softmax score, Cosine similarity, and Entropy of the softmax distribution as baseline uncertainty measures. Some VLMs use Sigmoid activation rather than Softmax [53, 49, 56, 22, 45] – for these methods, we only report the Sigmoid score as its results are identical to the cosine similarity and the entropy is unsuitable as the scores are not normalised to a distribution. We do not impose a minimum score threshold on any predictions, as the AuPR metric captures performance over all possible thresholds.
@@ 4.4 Datasets
We focus on general object recognition and test with the large-scale and prevailing datasets used to benchmark image classification and object detection: For image classification, we test with the ImageNet1k validation dataset [9, 35]. It contains $50,000$ images from $1000$ classes and is publicly available for research. For object detection, we test with the MS COCO 2017 validation dataset [23]. It contains $4,952$ images with $36,781$ annotated objects from $80$ classes and is publicly available for research. We show results for domain-specific classification datasets in the Supplementary Sec. S4.
@@ 5 Experiments and Results
AuROC AuPR P@95R R@95P TP OSE Accuracy
Classifier Uncertainty $\uparrow$ $\uparrow$ $\uparrow$ $\uparrow$ $\uparrow$ $\downarrow$ $\uparrow$
CLIP [34] Softmax 79.7 71.5 46.2 3.4 31,023 50,000 62.1
Cosine 72.2 61.3 42.5 0.1
Entropy 80.2 72.3 45.4 2.8
ALIGN [19] Softmax 81.0 74.7 48.1 8.3 32,618 50,000 65.2
Cosine 72.4 62.1 44.8 -
Entropy 80.9 75.0 46.8 8.6
CoCa [47] Softmax 79.0 71.7 46.0 3.4 31,725 50,000 63.4
Cosine 73.2 63.0 43.3 0.1
Entropy 80.5 73.0 46.8 3.5
ImageBind [14] Softmax 82.8 79.1 52.8 5.0 38,405 50,000 76.8
Cosine 79.2 73.9 50.7 0.2
Entropy 84.3 80.1 56.3 5.6
SigLIP [53] Sigmoid 81.0 76.2 52.1 2.2 37,851 50,000 75.7
LanguageBind [59] Softmax 83.9 80.7 54.7 10.5 39,243 50,000 78.5
Cosine 81.6 77.4 52.8 0.8
Entropy 85.4 81.7 58.4 11.1
Tab. 1 shows that all six VLM classifiers perform poorly for open-set recognition. For scenarios that require high recall of true positive predictions (i.e. $95\%$ recall), approximately every second prediction returned is an open-set error (i.e. precision between $46.2\%$ and $58.4\%$). For scenarios that require high precision of predictions (i.e. $95\%$ precision), less than 12% of the true positive predictions are retained (i.e. recall between 3.4% and 11.1%). Comparing across uncertainty types, the VLM classifiers show better uncertainty performance when using the predicted Softmax score or Entropy rather than the predicted Cosine similarity. This suggests raw similarity in VLM feature space is not the best indicator of open-set error, with measures of relative similarity comparative to other classes offering better performance.
@@
ViLD (R152)[15] Softmax “background” 60.1 7.5 15.4 16,011 1,485,600 46.4
Cosine “background” 33.1 6.9 0.1
Entropy “background” 62.2 8.4 16.3
@@
High open-set error counts are not inherently a problem if they are produced with high uncertainty. Yet for all tested detectors, the baseline uncertainty methods cannot adequately distinguish true positive detections from open-set errors. When thresholding for high recall of true positives, at best only every fifth detection is a true positive (i.e. precision ranges between $5.9\%$ and $21.9\%$). When thresholding for high precision, at best less than half of true positives are recalled (i.e. recall ranges between 16.3% and 48.5%). We include some qualitative examples of highly confident open-set errors in the Supplementary Sec. S6.
Notably, the precision-recall characteristics for VLM classification and object detection differ significantly – VLM classifiers achieve better performance at high recall, whereas VLM detectors achieve better performance at high precision. This suggest that while the VLM classifiers struggle with over-confident open-set errors, the VLM detectors suffer from under-confident true positives. We include histograms that visualise this trend in the Supplementary Sec. S7.
We also consider how the performance of uncertainty differs for identifying open-set error versus closed-set misclassifications, with results and a detailed discussion in the Supplementary Sec. S5. We find that the performance of softmax uncertainty is lower for open-set error than for closed-set error in both classification and detection tasks. In classification, open-set performance is particularly decremented when thresholding for high precision – likely due to a greater proportion of overconfident open-set error – whereas detection primarily degrades with increased numbers of errors, with ViLD for example producing 11.7×11.7$\times$11.7 × more open-set than closed-set misclassifications.
@@ 5.2 Can Negative Queries Improve Open-set Performance?
Table 3: The open-set performance of state-of-the-art VLM classifiers on ImageNet1k when used with negative random words (R-W) or random embeddings (R-E).
@@
CLIP [34] 500 R-W 73.4 48.1 3.3 29,961 41,428 59.9
(Softmax) 2500 R-W 72.8 49.8 1.6 28,657 35,092 57.3
500 R-E 71.9 47.2 2.5 30,160 44,673 60.3
2500 R-E 71.7 47.1 0.9 29,291 40,497 58.6
@@
ALIGN [19] 500 R-W 75.3 48.9 10.4 31,731 43,198 63.5
(Softmax) 2500 R-W 74.6 49.5 8.9 30,588 37,738 61.2
500 R-E 74.9 48.7 8.6 32,342 47,228 64.7
2500 R-E 75.1 49.7 7.2 31,462 41,210 62.9
None 71.7 46.0 3.4 31,725 50,000 63.4
CoCa [47] 500 R-W 75.4 49.6 0.1 30,276 39,528 60.6
(Softmax) 2500 R-W 75.9 50.9 6.0 29,107 34,360 58.2
500 R-E 71.7 46.0 3.4 31,723 49,966 63.4
2500 R-E 71.7 46.1 3.4 31,719 49,866 63.4
@@
ImageBind [14] 500 R-W 83.8 63.1 7.7 35,859 29,172 71.7
(Softmax) 2500 R-W 84.6 66.1 8.3 34,179 22,872 68.4
500 R-E 81.0 57.3 9.0 36,966 38,636 73.9
2500 R-E 81.7 59.0 11.2 35,954 33,349 71.9
None 76.2 52.1 5.2 37,851 50,000 75.7
SigLIP [53] 500 R-W 77.6 56.7 2.3 36,111 33,287 72.2
(Sigmoid) 2500 R-W 78.8 60.2 2.4 34,507 26,110 69.0
500 R-E 84.7 73.0 3.6 22,772 9,570 45.5
2500 R-E 90.1 83.9 7.9 10,561 2,258 21.1
None 80.7 54.7 10.5 39,243 50,000 78.5
LanguageBind [59] 500 R-W 84.7 64.9 13.1 36,596 28,337 73.2
(Softmax) 2500 R-W 85.4 68.2 11.7 34,540 21,230 69.1
500 R-E 80.8 54.9 10.8 39,226 49,646 78.5
2500 R-E 80.9 55.1 10.9 39,210 49,032 78.4
Figure 2: The trade-off between closed-set and open-set performance with increasing number of negative queries for different VLM classifiers.
Classification: Negative class queries provide the VLM classifiers with the ability to select “unknown” when testing an input. When used in larger quantities, negative queries can capture and remove significant portions of open-set error, yet this happens in trade-off with a reduction in closed-set performance – see Tab. 3. For example, ImageBing using $2,500$ random word negative queries effectively halves the open-set error, but this occurs in trade-off with a closed-set accuracy reduction of $8.4\%$ (approximately $4,000$ of the closed-set images are incorrectly classified as one of the random words).
While negative queries noticeably reduce the raw open-set error count, this is not always mirrored with increased performance in the open-set uncertainty metrics. We present the change in closed-set and open-set performance against the number of negative query points in Fig. 2. Generally the use of negative random words appears the most effective, with CoCa, ImageBind, SigLIP and LanguageBind showing improved open-set performance with minimal loss of closed-set performance. For some of the classifiers, there is negligible change in open-set performance despite the reduction in open-set error. We explore this further in the Supplementary Sec. S8, where we show that in some VLMs the negative queries only capture and remove “easy” open-set errors, i.e. open-set errors that already have high uncertainty.
Object Detection: The VLM detectors do not show a consistent response to the use of negative queries. Different detectors show different sensitivities to negative query counts, different trade-off rates between closed-set and open-set performance, and differences in effectiveness of embedding versus word negative query types – see Tab. 4 and Fig. 3. Looking at Fig. 3, we can consider the sensitivity, trade-off rate and response of the detectors to negative queries: greater performance changes with increasing queries indicates sensitivity; large decreases in mAP with more queries indicates increased trade-off rates between open-set and closed-set performance; and a positive response to the negative queries presents as an increase in AuPR when increasing queries (negative response decreases AuPR).
@@
Figure 3: The trade-off between closed-set and open-set performance with increasing number of negative queries for different VLM object detectors.